# Class Agnostic Instance-level Descriptor for Visual Instance Search

### v1

@@ Abstract
Despite the great success of the deep features in content-based image retrieval, the visual instance search remains challenging due to the lack of effective instance level feature representation. Supervised or weakly supervised object detection methods are not among the options due to their poor performance on the unknown object categories. In this paper, based on the feature set output from self-supervised ViT, the instance level region discovery is modeled as detecting the compact feature subsets in a hierarchical fashion. The hierarchical decomposition results in a hierarchy of feature subsets. The non-leaf nodes and leaf nodes on the hierarchy correspond to the various instance regions in an image of different semantic scales. The hierarchical decomposition well addresses the problem of object embedding and occlusions, which are widely observed in the real scenarios. The features derived from the nodes on the hierarchy make up a comprehensive representation for the latent instances in the image. Our instance-level descriptor remains effective on both the known and unknown object categories. Empirical studies on three instance search benchmarks show that it outperforms state-of-the-art methods considerably.
Index Terms:
Instance search, Content-based Image Retrieval, Image search.
@@ I Introduction
Visual instance search is a more challenging search task compared to the conventional content-based image retrieval. It requires the search system to return not only the relevant images, in which the target instance is in presence, but also the localization of the target instance in the relevant images. The localization is usually given as the bounding boxes over the target instance regions. Since the instance only appears as an object in an image sub-region, it could be occluded or undergone various types of transformations. For this reason, features that only describe an image local such as SIFT [1], SURF [2], and DELF [3] are expected. However, although these features are able to represent an object local, a single local feature is usually too fine-grained to form a complete description about an object instance or meaningful part of it (e.g., a logo on a hat). Moreover, there is no way to integrate several local features to achieve that given no prior is provided. Due to the above reasons, the instance-level feature representation is preferred, which is expected to build feature description for a complete instance or a semantic part of an instance.
@@
However, it is non-trivial to design an instance-level feature that is suitable for the search task. First of all, the feature detector, which is responsible for the instance region localization, should be class-agnostic. Otherwise, the detector only performs well on the known object categories [4, 5]. For this reason, supervised or weakly supervised object classification/detection/segmentation models are out of consideration [6]. Moreover, the precise localization on the meaningful instance region or its sub-region is expected. On the one hand, it ensures the bounding boxes returned to the query are in high quality. Moreover, it also guarantees that the instance feature is derived from a precise instance region, which in turn impacts the search quality.
@@
Due to the aforementioned challenges, effective instance-level feature is still slow to occur. In [7], a cycle self-training framework is proposed for instance search, in which the instance search is addressed by an object tracking framework. The instance region is localized by the correlation operation between the query instance and the candidate image as the GlobalTrack [8]. Although good performance is achieved on several benchmarks, such kind of framework is hardly scalable to large-scale search tasks, where the scale of candidate images could be up to million or even billion level. In contrast to self-training framework, DASR (Deeply Activated Salient Region) [4] is able to extract semantically significant regions from the candidate images. The salient regions are obtained by back-propagating peaks on the class activation map of an image. Encouraging search performance is reported on two benchmarks. Unfortunately, the instance localization accuracy turns out to be low in particular for the unknown object categories as the localization in DASR is built upon a pre-trained classification ConvNet.
@@
Recently, the research in computer vision has been marked by the self-supervised ViT model [9]. The distinctive token-wise feature can be built after the self-distillation. These features are so semantically meaningful that they could emerge as the semantic segmentation on an input image. Since the ViT is trained without any label, it becomes the cornerstone for the unsupervised object/semantic segmentation. With the support of self-supervised ViT, each token on the output layer is modeled as a node in an affinity graph, the unsupervised segmentation task is therefore treated as a graph partition problem in [10, 11, 12]. In particular, TokenCut in [12] treats the segmentation as a bipartite graph partition problem. One sub-graph is treated as the foreground object and another is treated as the background. While CutLER in [10] cuts the graph in a chain. Without referring to any other images or labels, significantly better segmentation accuracy than the previous methods is achieved by CutLER and TokenCut.
@@
In this paper, the instance-level regions are detected by a hierarchical decomposition over patch-level features from self-supervised ViT. The decomposition is carried out by repeated bisecting clustering as shown in Figure 1. This process essentially segments one image repeatedly into semantically salient regions. Different from CutLER and TokenCut, we do not simply treat a segmented sub-region as background or a foreground object. As long as they are sufficiently compact, the segmented regions are good candidates of instance-level descriptors. For this reason, the repeated bisecting decomposition continues on the resulting sub-regions until the carefully designed termination condition is reached. The hierarchical decomposition well addresses the problem of object occlusion and object embedding, which are widely observed in the real scenarios.
In summary, the contributions of this paper are threefold. 1) we model the problem of instance-level feature detection as a hierarchical decomposition on an image. The decomposition is addressed by stochastic optimization. This leads to an efficient and relatively exhaustive description about the instances/sub-instances latent in an image. 2) “dummy node” and “terminate condition” are proposed to support the design of this instance-level feature detector. This detector is able to detect instances in different granularity and it is class-agnostic. 3) It shows superior performance on single instance search, multi-instance search as well as image search. This does indicate our method allows the image search and instance search to be unified under the same retrieval framework. To the best of our knowledge, this is the first feature could achieve so in the literature. We have made the codes of our work publicly available [^1] https://github.com/wlzhao22/CLAID.
@@ II-A Instance Search
Before the emergence of deep ConvNet, instance search is addressed as a sub-image retrieval task [13]. Image local features such as SIFT [1] and SURF [2] were widely adopted. Essentially, the search seeks for the correspondence of local features between the query instance and the candidate images. To speed-up the match, feature aggregation methods such as BoVW [14] and VLAD [15, 16] are introduced. Although much better efficiency is achieved, the matching of local features has been degraded to the measurement of similarity between the query instance and the whole candidate image. In addition, due to the sensitivity of local features to the object deformation and 3D rotation, such kind of search paradigm only works on rigid objects.
Due to the wide success of neural-networks in many multimedia tasks, deep features have been introduced to the instance search task [17, 18, 19, 20, 21, 22, 23] in the recent decade. Usually, an image is represented by a deep global feature, which is pooled from certain conv-layer. During the pooling, the higher weights are assigned to the potential instance regions. The feature turns out to be robust to various image transformations, whereas, lacks of distinctiveness. In addition, methods of this category are unable to localize the instance from the relevant images.
In the recent studies [24, 7], the similarity between the query instance and the candidate images is learned via a ConvNet. The correlation is performed between the query instance and each candidate image. High similarity is returned when the target instance appears in the image. Superior performance is reported on several benchmarks. However, the instance search is essentially performed via a deep convolution between the query and all the candidate images. It is hardly scalable to large-scale due to the high time costs.
Instead of borrowing features designed for other tasks, several attempts [25, 4] have been made to design instance-level feature for the search task. In general, there are two major steps involved in the methods, namely instance localization and feature extraction from the detected instance regions. Method in [25] is built upon Faster-RCNN [26], which limits its capability in detecting object from unknown categories. In contrast, DASR [4] relies only on a pre-trained image classification network. It shows much stable performance over [25] as it remains effective on the unknown object categories. However, the localization is imprecise as the detection fully relies on the class activation map. Although encouraging performance is achieved for method in [27], its performance is still unsatisfactory as its region detector relies on weakly supervised learning.
@@ II-B Unsupervised Semantic/Object Segmentation
Owing to the success of self-distilled ViT DINO [9], the high quality of unsupervised semantic/object segmentation/detection is achievable in the recent studies [10, 11, 12]. In these works, the segmentation task is fulfilled by decomposing an affinity matrix, which is built based on the patch-level features derived from DINO. In [11], spectral clustering is adopted for the decomposition. The major eigenvectors after the decomposition correspond to the foreground object in the image. In TokenCut [12] and CutLER [10], the affinity matrix is transformed into a graph by filtering out edges with low similarity. Thereafter, NCut [28] is adopted for the graph partition in both methods. The difference is that only one cut is applied on the graph in TokenCut. The image region that corresponds to the sub-graph with high energy is viewed as an object. While the other is treated as the background. CutLER applies TokenCut three times. The resulting background of each round is cut into foreground object and the background in the next round. Compared to TokenCut, considerably higher segmentation accuracy is achieved. The coarse segmentation formed by the decomposition is further refined by pairwise CRF [11] or MaskCut [10].
@@
Intuitively, methods in unsupervised semantic/object segmentation can be directly borrowed as the detector for the instance search. However, these two tasks are still essentially different. First of all, one cannot assume that the submitted instance query is a complete object. The visual query could be a meaningful sub-region of an object, i.e. a logo on the coat or a cap of the bottle. Or on another extreme, the query is comprised by several visual objects, which is known as multiple-instance search. Theoretically speaking, any semantic part of an image could be cropped out as the query. The instance level feature detector should detect out these semantic regions from one image as much as possible. Different from TokenCut [12] and CutLER [10], a novel hierarchical decomposition procedure is designed in our paper. Given an image is represented by a collection of patch features output from a ViT layer, one feature corresponds to a patch on the ViT input layer. A bisecting procedure is proposed to partition the feature set repeatedly into a bunch of subsets. Each subset corresponds to a semantic region in the image. Instead of simply fixing the number of layers of such bisecting, a condition is designed to judge whether the bisecting should terminate at certain subset.
@@
As a result, similar as the conventional local feature, the collection of such features makes up a full description for the whole image. Whereas, unlike the conventional local feature, the descriptions are instance-level and semantically meaningful. More importantly, in contrast to conventional local features, typically only around 30 descriptors are extracted from one image, which largely relieves the burden of afterwards indexing and nearest-neighbor search.
@@ III-A Feature Detection by Hierarchical Decomposition
In ViT [29], an image is partitioned into small patches in equal size like a jigsaw puzzle, an instance inside the image is composed by dozens of such small patches. When feeding into the backbone, one patch is represented into a feature vector. Given $X$ is the set of features produced by the backbone, an instance inside the image corresponds to a subset of features. According to  [9], the pairwise feature distances within the subset are expected to be small, while their distances to the features outside the subset should be relatively larger. Therefore, detecting semantically salient regions could be modeled as a problem of clustering that minimizes the intra-cluster distances. Given there are k clusters $\{S_{1},\cdots,S_{k}\}$, our objective is
[…]
Although the objective looks similar to that of traditional k-means for the first sight, it is essentially different from traditional k-means and most of its variants. The objective minimizes the summation of pairwise $l_{2}$-distance within each cluster. Intuitively, we have to try out all the possible combinations of the feature vectors across different clusters to seek for the optimal solution for Eqn. 1, which is however computationally prohibitive. Moreover, the egg-chicken loop in the traditional k-means is no longer applicable as no cluster centroid is defined in Eqn. 1.
In our paper, the procedure proposed in k-sums [30] is employed to address this optimization problem. Given $x_{i}\in S_{w}$, the overall distance between sample $x_{i}$ and cluster $S_{w}$ is defined as
$d(x_{i},S_{w})=\sum_{x_{j}\in S_{w}}\parallel x_{i}-x_{j}\parallel^{2},$ (2)
which is the summation of distances between feature $x_{i}$ and the rest of features in $S_{w}$. Eqn. 2 can be further simplified as
@@
where $D_{w}=\sum_{x_{j}\in S_{w}}x_{j}$ and $n_{w}=|S_{w}|$. The first term and the last term in Eqn. 3 can be calculated conveniently given the $\textit{l}_{2}$-norms of each feature is pre-computed. The second term is the inner-product between feature $x_{i}$ and $D_{w}$, which is comparable to measuring the distance between $x_{i}$ and a centroid in the k-means. Since features are $l_{2}$-normalized in our problem, Eqn. 3 is simplified further as $d(x_{i},S_{w})=2{\cdot}n_{w}-2{\cdot}x_{i}^{\prime}{\cdot}D_{w}$. Given currently $x_{i}\in S_{w}$, we want to check whether Eqn. 1 would decrease assuming $x_{i}$ was moved to $S_{v}$ already,
@@
If $d(x_{i},S_{v})<d(x_{i},S_{w})$, such movement is favored. In particular, we can check over $k-1$ clusters to seek for the most greedy movement that Eqn. 1 decreases the most. In one round, each feature $x_{i}$ is selected by random to seek for the most greedy movement. While if $d(x_{i},S_{v})\geq d(x_{i},S_{w})$, the movement will not be undertaken. The greedy optimization usually converges to a local optima [30]. Since the step that seeks the best movement is comparable to seeking the closest centroid in traditional k-means, the computational complexity of above optimization is on the same par as the traditional k-means.
Repeated Bisecting The primary goal of clustering is to distinguish different instance regions. However, it is not appropriate to segment the image directly into k regions. First of all, the hyper-parameter k is hard to be set. The number of potential instance regions varies from one image to another. Moreover, our perception on visual instances is in multiple scales. For instance, the logo on the bottle is an instance. However, it is only a part of another instance when we view the bottle as an instance. Furthermore, when the bottle is on a desk, the bottle is just a small instance embedded in another bigger instance “desk”. In order to mimic the human perception about the visual instances, the clustering is carried out hierarchically. Specifically, the feature set is bisected into two each time. For a complete image as input, each step we decompose the image or its sub-image into two fine-grained sub-regions. Different levels of decomposition locate visual instances of different scales. A termination condition, which will be detailed later in this section, is designed to decide whether the bisecting should not be carried out further on a subset. By this way, we avoid setting the hyper-parameter k explicitly. The number of clusters we finally have is largely related to the number of latent instances in an image.
@@
In the following, the details are presented about how the hierarchical detection procedure is initialized and what is the condition to terminate the bisecting on a node. It is possible that the segmented regions are the empty background or the loose structures. A condition is further designed to filter out these dummy nodes on the hierarchy.

### v2

# Class Agnostic Instance-level Descriptor for Instance Search

@@ Abstract
Despite the great success of the deep features in content-based image retrieval, the visual instance search remains challenging due to the lack of effective instance-level feature representation. Supervised or weakly supervised object detection methods are not the appropriate solutions due to their poor performance on the unknown object categories. In this paper, based on the feature set output from self-supervised ViT, the instance-level region discovery is modeled as detecting the compact feature subsets in a hierarchical fashion. The hierarchical decomposition results in a hierarchy of instance regions. On the one hand, this kind of hierarchical decomposition well addresses the problem of object embedding and occlusions, which are widely observed in real scenarios. On the other hand, the non-leaf nodes and leaf nodes on the hierarchy correspond to the instance regions in different granularities within an image. Therefore, features in uniform length are produced for these instance regions, which may cover across a dominant image region, an integral of multiple instances, or various individual instances. Such a collection of features allows us to unify the image retrieval, multi-instance search, and instance search into one framework. The empirical studies on three benchmarks show that such an instance-level descriptor remains effective on both the known and unknown object categories. Moreover, the superior performance is achieved on single-instance and multi-instance search, as well as image retrieval tasks.
@@ I Introduction
Visual instance search is a more challenging search task compared to the conventional content-based image retrieval. It requires the search system to return not only the relevant images in which the target instance is present, but also the localization of the target instance in the relevant images. The localization is usually given as the bounding boxes over the target instance regions. Since the instance only appears as an object in an image sub-region, it could be occluded or undergo various types of transformations. For this reason, features that only describe an image local, such as SIFT [1], SURF [2], and DELF [3] are expected. However, although these features are able to represent an object local, a single local feature is usually too fine-grained to form a complete description about an object instance or meaningful part of it (e.g., a logo on a hat). Moreover, these features are vulnerable to object deformations and big viewpoint variations. Due to the above reasons, robust instance-level feature representation, which is able to describe a complete instance or a semantic part of an instance, is expected.
@@
However, it is non-trivial to design an instance-level feature that is suitable for the search task. First of all, the feature detector, which is responsible for the instance region localization, should be class-agnostic. Otherwise, the detector only performs well on the known object categories [4, 5]. For this reason, supervised or weakly supervised object classification/detection/segmentation models are out of consideration [6, 7]. Moreover, the localization of the instance region or its sub-region of various categories is expected. On the one hand, it ensures the target instance returned to the query are precisely localized in the reference images. On the other hand, it enables the instance feature to be derived from a precise instance region, which in turn guarantees the search quality.
@@
Due to the aforementioned challenges, an effective instance-level feature is still slow to occur. In [8], a cycle self-training framework is proposed for instance search, in which the instance search is addressed by an object tracking framework. The instance region is localized by the correlation operation between the query instance and the candidate image as the GlobalTrack [9]. Although good performance is achieved on several benchmarks, such a framework is hardly scalable to large-scale search tasks, where the scale of candidate images could be up to a million or even a billion. In contrast to the self-training framework, DASR (Deeply Activated Salient Region) [4] is able to extract semantically significant regions from the candidate images. The salient regions are obtained by back-propagating peaks on the class activation map of an image. Encouraging search performance is reported on two benchmarks. Unfortunately, the instance localization accuracy turns out to be low, in particular for the unknown object categories, as the localization in DASR is built upon a pre-trained classification ConvNet.
@@
Recently, the research in computer vision has been marked by the self-supervised ViT model [10]. The distinctive token-wise feature can be built after the self-distillation. These features are semantically meaningful. They could emerge as the semantic segmentation of an input image. Since the ViT is trained without any labels, it becomes the cornerstone for the unsupervised object/semantic segmentation. With the support of self-supervised ViT, each token on the output layer, which also corresponds to a fixed-size patch on the input image, is modeled as a node in an affinity graph. The unsupervised segmentation task is therefore treated as a graph partition problem in [11, 12, 13]. The decomposed subgraphs correspond to a number of subsets of patches on the input image. Each group of patches constitutes an object region. In particular, TokenCut in [13] treats the segmentation as a bipartite graph partition problem. One sub-graph is treated as the foreground object, and another is treated as the background. While CutLER in [11] cuts the graph in a chain. Without referring to any other images or labels, significantly better segmentation accuracy than the previous methods is achieved by CutLER and TokenCut.
@@
In this paper, the instance-level regions are detected by a hierarchical decomposition over the token-wise features from self-supervised ViT. The decomposition is carried out by repeated bisecting clustering as shown in Figure 1. The repeated bisecting decomposition continues on the resulting sub-regions until the carefully designed termination condition is reached. This process essentially segments one image repeatedly into semantically salient regions. Different from CutLER and TokenCut, we do not simply treat a segmented sub-region as background or a foreground object. As long as it is sufficiently salient, a feature is produced for the decomposed region. With such whole-to-part decomposition, one image is represented by a collection of features that are able to describe the image regions in multiple granularities. On the one hand, such decomposition well addresses the problem of object occlusion and object embedding, which are widely observed in real scenarios. On the other hand, it also builds the feature representation to support image retrieval, multi-instance search, and single instance search. In summary, the contributions of this paper are threefold.
  • We model the problem of instance-level feature detection as a hierarchical decomposition on the patch-level feature set of an image. The decomposition is addressed by stochastic optimization. This leads to an efficient and relatively exhaustive description of the instances/sub-instances latent in an image.
  • The “terminate condition” and “dummy node” are proposed to support the hierarchical decomposition. The former allows the decomposition to adapt to instances in different sizes and shapes and avoids a sensitive parameter, namely the depth of hierarchical decomposition. The latter helps to filter out many invalid features.
  • The hierarchical decomposition leads to a collection of feature representations for the latent instances in different granularities. They may represent a region large enough to support the similarity judgment between images. They may also represent a region composed of multiple instances or a single instance region. Therefore, our method allows the image search and instance search to be unified under the same retrieval framework.
To the best of our knowledge, this is the first feature that is able to fulfill three search tasks in the literature. We have made the codes of our work publicly available [^1] https://github.com/wlzhao22/CLAID.
@@ II-A Instance Search
Before the emergence of deep ConvNet, instance search was addressed as a sub-image retrieval task [14]. Image local features such as SIFT [1] and SURF [2] were widely adopted. Essentially, the search seeks the correspondence of local features between the query instance and the candidate images. To speed up the match, feature aggregation methods such as BoVW [15] and VLAD [16, 17] are introduced. Although much better efficiency is achieved, the matching of local features has been degraded to the measurement of similarity between the query instance and the whole candidate image. In addition, due to the sensitivity of local features to the object deformation and 3D rotation, this kind of search paradigm only works on rigid objects.
Due to the wide success of neural networks in many tasks, deep features have been introduced to the instance search task [18, 19, 20, 21, 22, 23, 24] in the recent decade. Usually, an image is represented by a deep global feature, which is pooled from a certain conv-layer. During the pooling, the higher weights are assigned to the potential instance regions. The feature turns out to be robust to various image transformations, whereas it lacks distinctiveness. In addition, methods of this category are unable to localize the instance from the relevant images.
In the recent studies [25, 8], the similarity between the query instance and the candidate images is learned via a ConvNet. The correlation is performed between the query instance and each candidate image. High similarity is returned when the target instance appears in the image. Superior performance is reported on several benchmarks. However, the instance search is essentially performed via a deep convolution between the query and all the candidate images. It is hardly scalable to a large scale due to the high time costs.
Instead of borrowing features designed for other tasks, several attempts [26, 4] have been made to design instance-level features for the search task. In general, there are two major steps involved in the methods, namely instance localization and feature extraction from the detected instance regions. Method in [26] is built upon Faster-RCNN [27], which limits its capability in detecting objects from unknown categories. In contrast, DASR [4] relies only on a pre-trained image classification network. It shows much stable performance over [26] as it remains effective on the unknown object categories. However, the localization is imprecise as the detection fully relies on the class activation map. Although encouraging performance is achieved for the method in [28], its performance is still unsatisfactory as its region detector relies on weakly supervised learning.
@@ II-B Unsupervised Semantic/Object Segmentation
Owing to the success of self-distilled ViT DINO [10], the high quality of unsupervised semantic/object segmentation/detection is achievable in the recent studies [11, 12, 13, 29, 30, 31]. In these works, the segmentation task is fulfilled by decomposing an affinity matrix, which is built based on the patch-level features derived from DINO. In [12], spectral clustering is adopted for the decomposition. The major eigenvectors after the decomposition correspond to the foreground object in the image. In TokenCut [13] and CutLER [11], the affinity matrix is transformed into a graph by filtering out edges with low similarity. Thereafter, NCut [32] is adopted for the graph partition in both methods. The difference is that only one cut is applied to the graph in TokenCut. The image region that corresponds to the sub-graph with high energy is viewed as an object. While the other is treated as the background. CutLER employs a method called MaskCut [11], in which NCut [32] is applied three times. In each round, the resulting background is further divided into a foreground object and a background for the next iteration. It yields considerably higher segmentation accuracy than TokenCut. The coarse masks produced by this decomposition are subsequently refined with a pairwise CRF [12]. MaskCut [11] and NCut [32] are adopted in U2Seg and VoteCut, respectively, to cut an image into semantic regions. In VoteCut [31], an image is cut into several regions based on the feature set produced by six ViT backbones, respectively. The final segmentation is fused by pixel-wise voting from the clustered regions. Compared to CutLER, a more precise semantic segmentation of the major objects in one image can be produced.
@@
Intuitively, methods in unsupervised semantic/object segmentation can be directly borrowed as the detector for the instance search. However, these two tasks are still essentially different. First of all, one cannot assume that the submitted instance query is a complete object. The visual query could be a meaningful sub-region of an object, i.e. a logo on the coat or a cap of the bottle. Or on another extreme, the query is comprised by several visual objects, which is known as multiple-instance search. Theoretically speaking, any semantic part of an image could be cropped out as the query. The instance-level feature detector should detect these semantic regions from one image as much as possible.
@@
Different from TokenCut [13] and CutLER [11], a novel hierarchical decomposition procedure is designed in our paper. Given that an image is represented by a collection of patch features output from a ViT layer, one feature corresponds to a patch on the ViT input layer. A bisecting procedure is proposed to partition the feature set repeatedly into a bunch of subsets. Each subset corresponds to a semantic region in the image. Instead of simply fixing the number of layers of such bisecting, a condition is designed to judge whether the bisecting should terminate at a certain subset. As a result, similar to the conventional local feature, the collection of such features makes up a full description for the whole image. Whereas, unlike the conventional local feature, the descriptions are instance-level and semantically meaningful. More importantly, in contrast to conventional local features, typically only around 30 descriptors are extracted from one image, which largely relieves the burden of afterwards indexing and nearest-neighbor search.
@@ III-A Feature Detection by Hierarchical Decomposition
In ViT [33], an image is partitioned into small patches of equal size like a jigsaw puzzle, and an instance inside the image is composed of dozens of such small patches. When feeding into the backbone, one patch is represented as a feature vector. Given $X$ is the set of features produced by the backbone, an instance inside the image corresponds to a subset of features. According to  [10], the pairwise feature distances within the subset are expected to be small, while their distances to the features outside the subset should be relatively larger. Therefore, detecting semantically salient regions could be modeled as a problem of clustering that minimizes the intra-cluster distances. Given there are k clusters $\{S_{1},\cdots,S_{k}\}$, our objective is
$\mbox{Min. }\sum_{r=1}^{k}\sum_{i,j\in S_{r}\&i<j}\parallel x_{i}-x_{j}\parallel^{2}.$ (1)
Although the objective looks similar to that of traditional k-means at first sight, it is essentially different from traditional k-means and most of its variants. The objective minimizes the summation of pairwise $l_{2}$-distance within each cluster. Intuitively, we have to try out all the possible combinations of the feature vectors across different clusters to seek the optimal solution for Eqn. 1, which is, however, computationally prohibitive. Moreover, the egg-chicken loop in the traditional k-means is no longer applicable as no cluster centroid is defined in Eqn. 1.
In our paper, the procedure proposed in k-sums [34] is employed to address this optimization problem. Given $x_{i}\in S_{w}$, the aggregated distance between sample $x_{i}$ and the rest of feature points in $S_{w}$ is defined as
$d(x_{i},S_{w})=\sum_{x_{j}\in S_{w}}\parallel x_{i}-x_{j}\parallel^{2}.$ (2)
Eqn. 2 can be further simplified into following form
@@
where $D_{w}=\sum_{x_{j}\in S_{w}}x_{j}$ and $n_{w}=|S_{w}|$. The first term and the last term in Eqn. 3 are both the $\textit{l}_{2}$-norm of the feature vectors. The second term is the inner-product between feature $x_{i}$ and $D_{w}$. Since features are $l_{2}$-normalized in our problem, Eqn. 3 is simplified further as $d(x_{i},S_{w})=2{\cdot}n_{w}-2{\cdot}x_{i}^{\prime}{\cdot}D_{w}$. The minimization on Eqn. 1 is fulfilled in a stochastic and greedy manner. Each time a feature $x_{i}\in S_{w}$ is randomly selected, we want to check whether Eqn. 1 would decrease assuming that $x_{i}$ was moved to $S_{v}$ already,
@@
If $d(x_{i},S_{v})<d(x_{i},S_{w})$, such movement would lead to the decrease of objective Eqn. 1. In particular, we can check over k1k-1 clusters to seek the most greedy movement that Eqn. 1 decreases the most. In one round, each feature $x_{i}$ is selected at random to seek the most greedy movement. While if $d(x_{i},S_{v})\geq d(x_{i},S_{w})$, the movement will not be undertaken. The greedy optimization usually converges to a local optima [34]. Since the step that seeks the best movement is comparable to seeking the closest centroid in traditional k-means, the computational complexity of the above optimization is on the same par as the traditional k-means.
Repeated Bisecting The primary goal of clustering is to distinguish different instance regions. However, it is not appropriate to segment the image directly into k regions. First of all, the hyperparameter k is hard to set. The number of potential instance regions varies from one image to another. Moreover, our perception of visual instances is on multiple scales. For instance, the logo on the bottle is an instance. However, it is only a part of another instance when we view the bottle as an instance. Furthermore, when the bottle is on a desk, the bottle is just a small instance embedded in another bigger instance “desk”. In order to mimic the human perception of the visual instances, the clustering is carried out hierarchically. Specifically, the feature set is bisected into two each time. For a complete image as input, in each step, we decompose the image or its sub-image into two fine-grained sub-regions. Different levels of decomposition locate visual instances of different scales. A termination condition, which will be detailed later in this section, is designed to decide whether the bisecting should not be carried out further on a subset. In this way, we avoid setting the hyper-parameter k explicitly. The number of clusters we finally have is largely related to the number of latent instances in an image.
@@
In the following, the details are presented about how the hierarchical detection procedure is initialized and what the condition is to terminate the bisecting on a node. It is possible that the segmented regions are the empty background or a small instance region embedded in a big empty background. It is unnecessary to build a feature for such regions. A condition is further designed to identify these nodes on the hierarchy.